OpenStudy (anonymous):

Solve equation: Cot3x=1

OpenStudy (anonymous):

cot3x=1 cos3x/sin3x=1, sin3x is not 0, 3x mustn't be k*pi, k is integer cos3x=sin3x when you multiply by sin3x This is possible only if 3x = pi/4 + k*pi, implies x=pi/12+k*pi/3